A solution contains 30% water and 70% milk. What quantity of water should be added to 6 litres of solution so that the water content became 40% ?
Answer by fractalier(6550)

Call x the amount of water you add...then we have
x(1) + 6(.30) = (6+x)(.40)
x + 1.8 = 2.4 + .4x
.6x = .6
x = 1 liter